A bag contains 5 red balls, numbered one through five, and 6 blue balls, numbered one through six. What is the probability that a ball chosen in random is a one or blue?
Not certain how to begin. :/

Answer by stanbon(75887)

A bag contains 5 red balls, numbered one through five, and 6 blue balls, numbered one through six. What is the probability that a ball chosen in random is a one or blue?
-------
P(one or blue) = P(one) + P(blue) - P(one AND blue)
----------
P(one or blue) = 2/11 + 6/11 - 1/11 = 7/11
==============================================
Cheer
Stan H.
=============